# Clutch Control: An Attention-based Combinatorial Bandit for Efficient Mutation in JavaScript Engine Fuzzing

### v1

@@ 6.1 Experimental Setup
At its core, Clutch is a contextual bandit applied to select locations in variable length (volatile) JavaScript test cases to fuzz JavaScript engines. We evaluate its effectiveness by inserting Clutch into three different state-of-the-art JavaScript engine fuzzers: Fuzzilli [20], FuzzJIT [45], and JIT-Picker [46]. We refer to these variants as FClutch, FJClutch, and JPClutch respectively. Each fuzzer applies the default number of mutations simultaneously, thus Clutch chooses arms at each timestep. We did not include OptFuzz [21], another extension of Fuzzilli, in the comparison as the available prototype relied on a version of llvm incompatible with current operating systems and JavaScript engines.
@@
To assess the performance of Clutch we compare with the particular view to efficiency. We consider the correctness of the test cases, i.e. the test cases that are valid JavaScript, not causing one of the four error types of JavaScript: SyntaxError, ReferenceError, TypeError, and RangeError. Generating JavaScript that does not cause such errors results in test cases that pass initial semantic and syntactic checks; in turn facilitating the deeper testing into the less frequently used paths in the engine. A common metric to test the ability of a fuzzer to reach these paths is code coverage. To measure the efficiency of Clutch we use the branch coverage per test case. Branch coverage is often used in the literature, providing a point of comparison with prior work [20].
@@
Each trial (JavaScript engine and fuzzer) is run for 24 hours and repeated 5 times, overall constituting 3,240 CPU hours across all experiments. Repeating experiments reduces the effect of randomness and allows for the statistical testing of different fuzzing approaches, as advocated for by Schloegel et al. [54]. In particular, we use the Mann-Whitney U test to determine statistical significance, and Vargha and Delaney’s test to quantify the magnitude of difference or effect size. The performance is considered significant when , and outperforms with a large effect size when .
@@
JIT-Picker and JPClutch achieve the highest rate of correctness of baseline fuzzers, and Clutch variants, respectively. JIT-Picker focuses on instrumentation and probing the engine using a subset of mutators from prior work [20], which causes this improvement in performance compared to the other approaches.
@@
Recent work from Caturano et al. [55] has argued that ‘more intelligent’ tools achieve the same (or better) results with a given number of attempts. As such, we report the efficiency as coverage per test case in Table 1. This allows us to determine how efficient each approach is with its test case budget. We can see that Clutch is generally able to improve the number of branches covered per execution, by an average of 8.9%. Indeed, Clutch increases the branches covered per execution consistently, doing so in six of the nine combinations of fuzzer and JavaScript engine. Specifically, each Clutch variant improves over the baseline fuzzer in two of the three engines. When comparing per JavaScript engine we see that Clutch always increases efficiency when fuzzing V8. This is due to a combination of the size of the engine and the resulting throughput of the fuzzer approaches. As the size of V8 is larger compared to JavaScriptCore and SpiderMonkey, the search space for Clutch is larger, giving it an advantage to target it efficiently. The throughput on V8 is also similar between Clutch variants and baseline fuzzers in comparison to the other JavaScript engines, which also accounts for efficiency improvements of Clutch, as baseline fuzzers achieve less coverage per test case.
@@ 7.3 Frequency of Branch Coverage
We have argued in favour of using Clutch to increase the efficiency of testing JavaScript engines. To demonstrate this further we compare branch frequency from the 10,000 generated test cases generated by Fuzzilli and FClutch on V8. We report these results in Figure 3. We can see that both Fuzzilli and FClutch find a comparable number of branches, covering 0.25%. of the V8 engine, accounting for over 34,000 branches. Comparable performance is also seen from the highest frequency of a single branch, 1320 for Fuzzilli and 1228 for FClutch. However, the frequency density of the the two highlights the difference in testing strategy. On average Fuzzilli requires 150.28 test cases to cover each branch, 23.5% more test cases than Clutch. Test cases from Fuzzilli trigger the same shallow paths due to higher proportion of errors they contain. Indeed, such results show that Clutch explores a maximal number of low frequency paths [11]. Clutch does so by learning to perform mutations in locations which are, a) valid for JavaScript, b) able to probe deeper in the engine more consistently.
@@ 8 Combinatorial Bandit Evaluation
We must also consider the performance of Clutch in relation to alternative bandit approaches. However, to do so in the context of JavaScript engines is challenging as it is computationally expensive to train bandits, and there is no ground truth to determine regret performance. Instead, we use evaluation methodologies used in prior work, providing a ground truth reward, and relatively inexpensive training time. Specifically, we consider the Gowalla Dataset in the same configuration as Nika et al. [10] for Volatile Combinatorial Bandits (VCBs), and use three hidden rewards as Hwang et al. [9] for CB. We use the hyperparameters and experimental setup as in the prior work, running experiments for 10,000 rounds and repeating 10 times. We present the results in Table 2.
@@
In this setting we consider a crowd sourcing problem used by Nika et al. [10], using the Gowalla dataset. At each round a random number of user check-ins are sampled, where each check-in is an arm. Context is the location (normalised longitude and latitude lying in ), and the reward is the ‘willingness to work’, i.e. battery status. Bandits must then select locations. We compare against two state-of-the-art bandits that handle the volatile nature of the arms: ACC-UCB [10] and CC-UCB [56]. From Table 2 we see that bandits improve when selecting an increasing number of arms due increased information gathered from the additional arms. Furthermore, Clutch has the best performance in both , achieving at least 78.1% fewer regret than other bandits.
@@
In this setting we consider three different, unknown, reward functions: , , and . Where is the set of contexts for arms and has the same dimension as and is randomly generated from a unit ball, remaining fixed during the experiment. Bandits must select from arms at each round, observing a context of features. Such an evaluation is inline with Hwang et al. [9] and Zhou et al. [7]. We compare against contextual neural bandits CN-UCB and CN-TS, and their linear equivalents CombLinUCB and CombLinTS [9]. We additionally include CN-TS(M=1), a special case of CN-TS drawing only one sample per arm. Volatile bandits require context features to be , which not originally considered by Hwang et al. [9] which uses 80 features (). For completeness we consider in Appendix D. From Table 2 we again see that the deep CBs outperform the linear counterparts. We also see that across all rewards Clutch outperforms the deep CBs, achieving between 4.1%-67.3% fewer regret.
@@ 9 Threats to validity
The selection of hyperparameters in a learning architecture may depend on domain knowledge, leading to bias, and may in general affect the performance of the model unpredictably. To mitigate this, Clutch controls its own rate of exportation, uses hyperparameters from prior work where applicable [49, 57, 48], and uses a grid-search to find optimal values for remaining hyperparameters (Appendix A).
We did not place Clutch inside an exhaustive list of all JavaScript engine fuzzers. In particular, we have excluded OptFuzz [21] and Superion [41]. OptFuzz, while publicly available, requires recompiling the llvm library with custom hooks, and does not fully document the dependencies required to do so. Superion is used for fuzzing both XML and JavaScript engines, extending the AFL fuzzing model. As such we considered the approach and underlying system out-of-scope. Indeed, it is non-trivial to place Clutch inside new fuzzers, even those based on similar architecture. Instead, we selected three state-of-the-art fuzzers to evaluate the performance of Clutch. We argue these fuzzers demonstrate the generalisability of our approach, as they use different approaches and testing mutations, ranging from targeting the JIT to differential testing. We further test each fuzzer and Clutch variant on three separate and commonly used JavaScript engines. This removes bias from test subjects, showing the real world application of Clutch.
To mitigate for randomness in the different approaches we conducted large scale experiments. We repeated experiments in Section 6, and Appendix C.3 five times, accounting for 3,240 compute hours of experiments. In reporting our results we have followed guidance from Schlosser et al. [54], reporting metrics which provide deeper statistical comparison between different sets of data. Additionally, in Section 7, we use 10,000 test cases to compare between Clutch and the fuzzer baselines to provide an in-depth study on their behavioural differences.
@@ Appendix B Concrete Dropout
Dropout randomly samples from a masked weight matrix in the neural network [57]. This dropout probability and weight matrix can be approximated to the probabilistic deep Gaussian process [58]. By taking this approximation we can apply a Bayesian deep learning model to a neural network of layers and variations parameters. This allows us to place a prior on the network weights and then learn the posterior distribution over the weights () [49]. This can be formally written as the objective functions:
@@
The KL term can be considered the ‘regularisation’ term to ensure that the posterior and prior distributions do not diverge too far. Gal et al. [49] show that the KL term can be approximated to depend only on the probability of turning off a Bernoulli random variable (node) with with probability . From this we can derive Concrete Dropout [49], where we continuously relax dropout’s discrete mask.
@@ C.3 Comparison to ClutchLite
Thus we introduce ClutchLite, an alternative bandit model based on the Adaptive Contextual Combinatorial Upper Confidence Bound (ACC-UCB) [10] algorithm. While ACC-UCB was originally used for solving ‘dynamic resource allocation problems’ using a tree based structure, we have modified it to work within Clutch, by providing it with the ability to increase the search space over time (i.e. the number of instructions seen from test cases over time). ClutchLite can then handle the context of selecting instructions from the IR, and the variable number of arms that are seen over the course of fuzzing. As a result we can compare Clutch to another contextual bandit for selecting locations to mutate. We compare Clutch and ClutchLite for their efficiency of branches found per execution, reporting the results in Table 5.
@@ Appendix D Additional Combinatorial Bandit Evaluation
We include here the plots pertaining to the experiments detailed in Section 8 and Table 2. Figure 7 contains the regret of bandits in the volatile setting of the Gollawa dataset used by Nika et al. [10]. Table 6 includes the statistical testing values related to effect size (significant when ) and statistical significance (significant when ).

### v2

@@ 6.1 Experimental Setup
At its core, Clutch is a contextual bandit applied to select locations in variable length (volatile) JavaScript test cases to fuzz JavaScript engines. We evaluate its effectiveness by inserting Clutch into three different state-of-the-art JavaScript engine fuzzers: Fuzzilli [gros_fuzzilli_2023], FuzzJIT [wang_fuzzjit_nodate], and JIT-Picker [bernhard_jit-picking_2022]. We refer to these variants as FClutch, FJClutch, and JPClutch respectively. Each fuzzer applies the default number of mutations simultaneously, thus Clutch chooses arms at each timestep. We did not include OptFuzz [wang_optfuzz_2024], another extension of Fuzzilli, in the comparison as the available prototype relied on a version of llvm incompatible with current operating systems and JavaScript engines.
@@
To assess the performance of Clutch we compare with the particular view to efficiency. We consider the correctness of the test cases, i.e. the test cases that are valid JavaScript, not causing one of the four error types of JavaScript: SyntaxError, ReferenceError, TypeError, and RangeError. Generating JavaScript that does not cause such errors results in test cases that pass initial semantic and syntactic checks; in turn facilitating the deeper testing into the less frequently used paths in the engine. A common metric to test the ability of a fuzzer to reach these paths is code coverage. To measure the efficiency of Clutch we use the branch coverage per test case. Branch coverage is often used in the literature, providing a point of comparison with prior work [gros_fuzzilli_2023].
@@
Each trial (JavaScript engine and fuzzer) is run for 24 hours and repeated 5 times, overall constituting 3,240 CPU hours across all experiments. Repeating experiments reduces the effect of randomness and allows for the statistical testing of different fuzzing approaches, as advocated for by Schloegel et al. [schloegel_sok_2024]. In particular, we use the Mann-Whitney U test to determine statistical significance, and Vargha and Delaney’s test to quantify the magnitude of difference or effect size. The performance is considered significant when , and outperforms with a large effect size when .
@@
JIT-Picker and JPClutch achieve the highest rate of correctness of baseline fuzzers, and Clutch variants, respectively. JIT-Picker focuses on instrumentation and probing the engine using a subset of mutators from prior work [gros_fuzzilli_2023], which causes this improvement in performance compared to the other approaches.
@@
Recent work from Caturano et al. [caturano_discovering_2021] has argued that ‘more intelligent’ tools achieve the same (or better) results with a given number of attempts. As such, we report the efficiency as coverage per test case in Table 1. This allows us to determine how efficient each approach is with its test case budget. We can see that Clutch is generally able to improve the number of branches covered per execution, by an average of 8.9%. Indeed, Clutch increases the branches covered per execution consistently, doing so in six of the nine combinations of fuzzer and JavaScript engine. Specifically, each Clutch variant improves over the baseline fuzzer in two of the three engines. When comparing per JavaScript engine we see that Clutch always increases efficiency when fuzzing V8. This is due to a combination of the size of the engine and the resulting throughput of the fuzzer approaches. As the size of V8 is larger compared to JavaScriptCore and SpiderMonkey, the search space for Clutch is larger, giving it an advantage to target it efficiently. The throughput on V8 is also similar between Clutch variants and baseline fuzzers in comparison to the other JavaScript engines, which also accounts for efficiency improvements of Clutch, as baseline fuzzers achieve less coverage per test case.
@@ 7.3 Frequency of Branch Coverage
We have argued in favour of using Clutch to increase the efficiency of testing JavaScript engines. To demonstrate this further we compare branch frequency from the 10,000 generated test cases generated by Fuzzilli and FClutch on V8. We report these results in Figure 3. We can see that both Fuzzilli and FClutch find a comparable number of branches, covering 0.25%. of the V8 engine, accounting for over 34,000 branches. Comparable performance is also seen from the highest frequency of a single branch, 1320 for Fuzzilli and 1228 for FClutch. However, the frequency density of the the two highlights the difference in testing strategy. On average Fuzzilli requires 150.28 test cases to cover each branch, 23.5% more test cases than Clutch. Test cases from Fuzzilli trigger the same shallow paths due to higher proportion of errors they contain. Indeed, such results show that Clutch explores a maximal number of low frequency paths [bohme_coverage-based_2016]. Clutch does so by learning to perform mutations in locations which are, a) valid for JavaScript, b) able to probe deeper in the engine more consistently.
@@ 8 Combinatorial Bandit Evaluation
We must also consider the performance of Clutch in relation to alternative bandit approaches. However, to do so in the context of JavaScript engines is challenging as it is computationally expensive to train bandits, and there is no ground truth to determine regret performance. Instead, we use evaluation methodologies used in prior work, providing a ground truth reward, and relatively inexpensive training time. Specifically, we consider the Gowalla Dataset in the same configuration as Nika et al. [nika_contextual_2020] for Volatile Combinatorial Bandits (VCBs), and use three hidden rewards as Hwang et al. [hwang_combinatorial_2023] for CB. We use the hyperparameters and experimental setup as in the prior work, running experiments for 10,000 rounds and repeating 10 times. We present the results in Table 2.
@@
In this setting we consider a crowd sourcing problem used by Nika et al. [nika_contextual_2020], using the Gowalla dataset. At each round a random number of user check-ins are sampled, where each check-in is an arm. Context is the location (normalised longitude and latitude lying in ), and the reward is the ‘willingness to work’, i.e. battery status. Bandits must then select locations. We compare against two state-of-the-art bandits that handle the volatile nature of the arms: ACC-UCB [nika_contextual_2020] and CC-UCB [chen_contextual_2018]. From Table 2 we see that bandits improve when selecting an increasing number of arms due increased information gathered from the additional arms. Furthermore, Clutch has the best performance in both , achieving at least 78.1% fewer regret than other bandits.
@@
In this setting we consider three different, unknown, reward functions: , , and . Where is the set of contexts for arms and has the same dimension as and is randomly generated from a unit ball, remaining fixed during the experiment. Bandits must select from arms at each round, observing a context of features. Such an evaluation is inline with Hwang et al. [hwang_combinatorial_2023] and Zhou et al. [zhou_neural_2020]. We compare against contextual neural bandits CN-UCB and CN-TS, and their linear equivalents CombLinUCB and CombLinTS [hwang_combinatorial_2023]. We additionally include CN-TS(M=1), a special case of CN-TS drawing only one sample per arm. Volatile bandits require context features to be , which not originally considered by Hwang et al. [hwang_combinatorial_2023] which uses 80 features (). For completeness we consider in Appendix D. From Table 2 we again see that the deep CBs outperform the linear counterparts. We also see that across all rewards Clutch outperforms the deep CBs, achieving between 4.1%-67.3% fewer regret.
@@ 9 Threats to validity
The selection of hyperparameters in a learning architecture may depend on domain knowledge, leading to bias, and may in general affect the performance of the model unpredictably. To mitigate this, Clutch controls its own rate of exportation, uses hyperparameters from prior work where applicable [gal_concrete_2017, collier_deep_2018, vinyals_pointer_2017], and uses a grid-search to find optimal values for remaining hyperparameters (Appendix A).
We did not place Clutch inside an exhaustive list of all JavaScript engine fuzzers. In particular, we have excluded OptFuzz [wang_optfuzz_2024] and Superion [wang_superion_2019]. OptFuzz, while publicly available, requires recompiling the llvm library with custom hooks, and does not fully document the dependencies required to do so. Superion is used for fuzzing both XML and JavaScript engines, extending the AFL fuzzing model. As such we considered the approach and underlying system out-of-scope. Indeed, it is non-trivial to place Clutch inside new fuzzers, even those based on similar architecture. Instead, we selected three state-of-the-art fuzzers to evaluate the performance of Clutch. We argue these fuzzers demonstrate the generalisability of our approach, as they use different approaches and testing mutations, ranging from targeting the JIT to differential testing. We further test each fuzzer and Clutch variant on three separate and commonly used JavaScript engines. This removes bias from test subjects, showing the real world application of Clutch.
To mitigate for randomness in the different approaches we conducted large scale experiments. We repeated experiments in Section 6, and Appendix C.3 five times, accounting for 3,240 compute hours of experiments. In reporting our results we have followed guidance from Schlosser et al. [schloegel_sok_2024], reporting metrics which provide deeper statistical comparison between different sets of data. Additionally, in Section 7, we use 10,000 test cases to compare between Clutch and the fuzzer baselines to provide an in-depth study on their behavioural differences.
@@ Appendix B Concrete Dropout
Dropout randomly samples from a masked weight matrix in the neural network [collier_deep_2018]. This dropout probability and weight matrix can be approximated to the probabilistic deep Gaussian process [damianou_deep_2013]. By taking this approximation we can apply a Bayesian deep learning model to a neural network of layers and variations parameters. This allows us to place a prior on the network weights and then learn the posterior distribution over the weights () [gal_concrete_2017]. This can be formally written as the objective functions:
@@
The KL term can be considered the ‘regularisation’ term to ensure that the posterior and prior distributions do not diverge too far. Gal et al. [gal_concrete_2017] show that the KL term can be approximated to depend only on the probability of turning off a Bernoulli random variable (node) with with probability . From this we can derive Concrete Dropout [gal_concrete_2017], where we continuously relax dropout’s discrete mask.
@@ C.3 Comparison to ClutchLite
Thus we introduce ClutchLite, an alternative bandit model based on the Adaptive Contextual Combinatorial Upper Confidence Bound (ACC-UCB) [nika_contextual_2020] algorithm. While ACC-UCB was originally used for solving ‘dynamic resource allocation problems’ using a tree based structure, we have modified it to work within Clutch, by providing it with the ability to increase the search space over time (i.e. the number of instructions seen from test cases over time). ClutchLite can then handle the context of selecting instructions from the IR, and the variable number of arms that are seen over the course of fuzzing. As a result we can compare Clutch to another contextual bandit for selecting locations to mutate. We compare Clutch and ClutchLite for their efficiency of branches found per execution, reporting the results in Table 5.
@@ Appendix D Additional Combinatorial Bandit Evaluation
We include here the plots pertaining to the experiments detailed in Section 8 and Table 2. Figure 7 contains the regret of bandits in the volatile setting of the Gollawa dataset used by Nika et al. [nika_contextual_2020]. Table 6 includes the statistical testing values related to effect size (significant when ) and statistical significance (significant when ).